# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v2

@@ Results
We now turn ourselves towards the sources of scholarly attention (see Figs. 5 & 6). The contributions of different countries to the rise in foreign scholarly attention show significant geographical disparities (Fig. 5-A). Western nations, such as those in Western Europe, North America, and Australia—known for their high scholarly output—account for the largest share of this growth. The United States leads this growth, with an average increase of 172.3 additional papers per year compared to the pre-Arab Spring period, followed by the United Kingdom with 118.3 additional documents per year. Malaysia and Saudi Arabia also contribute significantly despite their lower scholarly output. In contrast, contributions from the Global South remain relatively limited.
[…]
@@
To understand how the Arab Spring impacted how foreign countries distributed their attention, we examine the change in positions of the target countries within the ranking of countries on which they focus scholarly attention. We show the change in ranks of target countries for the 10 most important sources of attention in Fig. S4 in the Supplementary Material. A one-way signed rank test indicates a statistically significant increase in the rankings of the government overthrow and civil war groups across all source countries—except for Germany, France, and the UK—at the 0.05 significance level. We find that Syria, Libya, Tunisia, and Jordan’s ranks generally increased in the post-Arab Spring period. Egypt’s rank was already relatively high before the Arab Spring and remained relatively stable. This suggests that the shift in scholarly attention toward these target countries is also significant from the perspective of the source countries. Such a trend may reflect an academic response to regional developments and possible changes in funding priorities or broader research policy shifts.
@@
The distribution of foreign attention’s contribution shares ($A_{i}^{*}$) follows a long-tailed pattern, making upward movement in the rankings more challenging at lower ranks due to the inertia created by established research infrastructures. For example, a country advancing from rank 2 to rank 1 requires significantly more publications than a country moving from rank 146 to rank 147 among the 147 countries worldwide in our work. To account for these disparities, the metric rank inertia  (?), $RI(A_{i}^{*})=\frac{\Delta r(A_{i}^{*})}{N-r(A_{i}^{*})}$, adjusts changes in contributions for country $i$. Here, $N$ represents the total number of countries (in this case, $N=147$) and $r(A_{i}^{*})$ denotes the ranking of the country $i$ within the global distribution of scholarly attention directed toward target countries. In Figure 5-B, we show that the countries exhibiting the greatest rank inertia are Malaysia (0.73), Saudi Arabia (0.61), the United Kingdom (0.54), Iran (0.54), and Qatar (0.48), where more rank inertia indicate larger changes in terms of relative ranks. The figure highlights the noticeable impact of the Arab Spring, particularly in countries with Arab or Muslim-majority populations and those within the region. Notably, Saudi Arabia and the United Kingdom are among the top five countries in absolute scholarly attention changes and rank inertia, highlighting their substantial influence in research related to the target countries.
@@
The pattern in Arab Spring-related research differs (see Fig. S4 in the Supplementary Materials). While the United States and the United Kingdom contributed significant scholarly attention, other Western countries such as France, Germany, Italy, Spain, and Canada were also important. Türkiye, Egypt, and Israel—countries within the region and directly or indirectly affected by the Arab Spring—also showed notable research activity. When examining rankings, Malaysia ranks 24th with 32 publications, while Saudi Arabia ranks 30th with only 23. This suggests that despite their broader research interests in the target countries, their focus does not extend significantly to the Arab Spring itself. A key reason for this may be that most Arab Spring-related research is concentrated in the social sciences, with limited exploration in other disciplines (Fig. S6 in Supplementary Materials).
@@
Figure 6: Repartition of scholarly attention towards target countries by discipline and source country: Differences in yearly scholarly attention by country, listing the top 10 countries with the greatest differences in the field of (A) Social sciences, (B) Health Sciences, (C) Life Sciences, and (D) Physical Sciences.
@@
Figures 6A–D show the top ten sources of foreign attention towards the region by computing the differences in yearly attention before and after the Arab Spring in the four main disciplines, thereby helping to understand how geographic, historical, and political factors shape research interests. Saudi Arabia has emerged as an important source of scholarly attention in the health, physical, and life sciences—mainly focusing on Egypt, followed by Jordan and Yemen—and exhibits less engagement in the social sciences. Within the same discipline, Iran has an interest in Oman. Meanwhile, Malaysia stands out as a key contributor in social and health sciences, though its attention remains primarily on Yemen. Spain concentrates largely on Morocco and secondarily on Tunisia, likely influenced by geographic proximity. Western countries—especially the United States, the UK, and France—show strong attention to social sciences, with Egypt and Syria receiving the most of their attention. Interestingly, the United States ranks only eighth in life sciences, while the UK and Malaysia are absent in the top 10 in this field.
@@
To assess the impact of the Arab Spring on patterns of scholarly attention, we employ a difference in differences (DiD) statistical framework  (?). This approach allows us to estimate the average effect of the Arab Spring by comparing changes over time in attention to countries directly affected by it, i.e., our target countries, or in the three subgroups (government overthrow, civil war, and governmental changes) with those observed in either the rest of the world or a subset of MENA countries that remained relatively unaffected during the same period (control groups). See the Materials and Methods section for the details.
@@
Log Domestic Attention log GDP, log Rpop. Rest of world
@@
Table 1: Difference in differences estimates of the effect of Arab Spring on scholarly attention. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of Arab Spring on log attention (total, domestic, foreign) and funding (total, domestic, foreign). Models control for log GDP, population, and research population where applicable. Each colored column represents estimates from separate regressions and compares the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes regressions where the prior trend parallel test fails to reject the null hypothesis of no difference in pre-Arab Spring trends at a 5% significance level.
Table S5 presents the difference-in-differences (DiD) analysis of scholarly attention—measured as the logarithm of total, domestic, and foreign attention—while controlling for log(GDP per capita), log(population), and log(researcher population). The analysis reveals that target countries affected by the Arab Spring experienced a statistically significant increase ($p$-value = 0.0129) in total scholarly attention compared to the rest of the world. At the subgroup level, countries with government overthrows and civil wars show a positive treatment effect on total attention, with the civil war subgroup exhibiting a stronger impact. In contrast, the effect for countries undergoing governmental changes is only statistically significant when they are compared to the rest of the MENA region, and they gained less attention relative to them. Further disaggregation by the source of attention offers additional insights. The domestic attention decreased significantly in the civil war subgroup compared to the rest of the world and the MENA region, whereas the government overthrow subgroup displays a positive effect—possibly reflecting reform strategies. Regarding foreign attention, both the government overthrow and civil war subgroups are linked to significant gains, with the civil war subgroup exhibiting a notably higher effect. Conversely, the governmental changes subgroup consistently records lower increases in foreign attention compared to rest of the world.
Collectively, these findings underscore the heterogeneous impacts of the Arab Spring on scholarly attention. While political instability via government overthrows and civil wars exhibited increased academic interest, governmental changes may dampen this engagement. These results highlight the possible importance of post-upheaval political trajectories in shaping scholarly discourse.
@@
Figure 7: Shifts in funding patterns before and after the Arab Spring. (A) Changes in funding before and after the Arab Spring per target country. The dashed line, dotted line, and shaded gray area represent the world’s mean, median, and interquartile range. We observe substantial increases for Egypt, Tunisia, Jordan, and Morocco. (B) & (C) Average domestic and foreign scholarly funding before and after the Arab Spring per target country. Domestic funding surged for Egypt and the governmental changes (GC) group, particularly Jordan and Kuwait. Foreign funding grows for the government overthrow group (GO), Morocco, and Jordan. (D) Total funding linked with scholarly attention toward the target countries per discipline. Plain and hashed bars indicate values before and after the Arab Spring, respectively. Although Social Sciences holds the smallest share of total funding, its allocation shows the biggest increases from 12.4% to 17.6%. (E) Change in scholarly attention towards the target countries as a function of changes in funding per source country. The strong correlation (Spearman rank correlation of 0.88) suggests that increases in funding may serve as a key explanatory factor for the observed rise in foreign scholarly engagement across target countries. Countries such as the United States, United Kingdom, France, Spain, and Saudi Arabia rank highest in both funding changes and scholarly attention.
@@
The increase in scholarly attention observed in target countries may correlate with other influential academic factors, e.g., funding. Since each researcher finances their research with allocated budgets, the topics of articles are influenced mainly by the goals of the received funding. We extract the countries providing funding from the article’s acknowledgment section and count funding in terms of the number of papers (see Materials and Method). Similarly to Scholarly attention, we model the funding patterns with a directed weighted network. Nodes once more represent countries, and directed edges capture the funding allocated from country $i$ to study country $j$, namely $F^{j}_{i}$ (see Materials and Method for details).
@@
Figure 7-A displays the changes in funding, $\Delta F_{i}^{*}$, received before and after the Arab Spring. The rise in funding for Egypt, Tunisia, Jordan, and Morocco significantly falls between the median and the third quartile, while the change for the rest is below the median (Syria, Kuwait, and Oman) or the first quartile (Libya, Yemen, and Bahrain). Domestic funding for studying Egypt and the governmental changes group, particularly Jordan and Kuwait, has significantly increased (Fig 7-B). The change for Kuwait was notable, from below the median to above the median. Bahrain experienced a decline. Foreign funding for research on government-overthrow groups, as well as Morocco and Jordan, has grown substantially (Fig 7-C).
Similarly to scholarly attention patterns, the total funding linked with scholarly attention toward the target countries, $F^{*}_{tot}$ rose for all disciplines after the Arab Spring (Fig. 7-D). Although the social sciences discipline received the smallest share of the total funding in both periods, its allocated share increased from 12.4% to 17.6%—a 5.2% uptick, suggesting that its growth in funding may partially explain the heightened social sciences scholarly attention.
[…]
@@
To assess whether the Arab Spring significantly influenced funding patterns, we applied the same difference-in-differences framework that we applied to the scholarly attention (refer to Materials and Methods), using the log of total, domestic, and foreign funding as outcome variables while controlling for the log of population and GDP per capita. The results, presented in Table S5 of the Supplementary Materials, indicate that the pre-treatment parallel trends assumption is often not satisfied, and the model fit is generally poor. Therefore, we refrain from interpreting these regression results in detail.
Nonetheless, the remaining analysis complements the patterns observed in Figure 7, offering a some statistical assessment of the Arab Spring’s impact on funding (refer to Table S5 in Supplementary Materials). The Civil War group shows a significant post-Arab Spring increase in total funding relative to the rest of the world ( p $<$ 0.01). In terms of funding sources, domestic support grew markedly for the governmental overthrown group (p $<$ 0.01). Foreign funding also rose significantly for the civil war group (p $<$ 0.01). Together, these results underscore a post-uprising shift in both the level and source of funding, varying by political trajectory.
@@
When a significant social or political phenomenon occurs, such as a revolution or civil war, a country’s institutions, including academia, are affected. One critical aspect influenced by these events is the mobility of scientists and researchers. After such an event, two main changes can occur regarding researcher mobility: the country may experience a brain drain, where the number of researchers leaving the country exceeds those entering, or researchers who had previously left due to political oppression and dissatisfaction with the former government may return as conditions improve. In such a situation, the net migration rate of researchers increases, and in the previous state, it decreases.
@@
In the context of increased scholarly attention, two scenarios can be considered. First, scientists who leave their country due to political upheavals may remain interested in their homeland and continue conducting related research. Second, scientists who return to their country after positive political and economic changes may be motivated to contribute to their country’s advancement. This leads to increased academic and collective knowledge through their intensified research efforts.
With this background, we investigate whether the rise in scholarly attention in target countries originated from all countries worldwide and corresponds with an increase in scholarly migration to those countries. Specifically, we seek to determine the rise in research interest from researchers in other countries, predominantly from those countries that have been destinations for migrating researchers from the target countries.
@@
Figure 8: Shifts in migration patterns before and after the Arab Spring (A) Net migration of researchers for target countries before and after the Arab Spring. In all panels, the gray area represents the interquartile range for global trends, with the dashed line indicating the median and the dotted line showing the mean. (B) Change in the number of migrations from the target countries to destination countries before and after the Arab Spring. We show the top 10 destinations. Islamic and Western countries are among the top 10 migration destinations for researchers from target countries, with special bonds between Saudi Arabia and Egypt. (D) Change in scholarly attention towards the target countries as a function of changes in migrations from the target countries.
We tracked scholarly migration from and to the target countries using the scholarly migration database created using changes in affiliations recorded in Scopus-indexed articles  (?, ?). Figure 8 shows the shifts in migration patterns before and after the Arab Spring.
@@
Before the Arab Spring, Egypt consistently attracted researchers, evidenced by positive net migration rates, except in 2008, when the rate nearly reached zero, which could be attributed to an increased outflow of researchers due to unfavorable economic conditions (Fig 8-A). With the onset of the Arab Spring, an increase in the net migration rate of researchers is observable. This may be due to two factors: researchers who had previously left the country because of economic inefficiencies and political turmoil might have returned due to improved conditions, and Egypt, being a relatively stable country in the region with reputable universities and a common language, could have attracted researchers from other target countries experiencing more adverse conditions. However, after 2015, the net migration rate tends to decrease, potentially due to the effects of the Arab Winter, referring to the resurgence of authoritarianism and Islamic extremism in some target countries. Tunisia exhibits a downward trend, transitioning from a country that attracted talent to one experiencing a brain drain. This process began in 2011, and the only subsequent year when the net migration rate was not negative was 2016—a significant political year in Tunisia. That year was politically substantial for Tunisia, marked by employment-related protests and the founding of the Democratic Party. Syria began experiencing a brain drain immediately following the Arab Spring. In contrast, data for Yemen and Libya are less informative due to their relatively small number of researchers compared to other countries. No clear pattern is observed within the group of countries undergoing governmental changes.
@@
We model scholarly migration as a directed network, $M$, where nodes represent countries and directed edges, $M_{i}^{j}$, indicate the number of scholars migrating from country $i$ to country $j$. Using a notation similar to that of funding and attention, the total emigration from the target countries to country $i$ is $M^{i}_{*}=\sum_{i\in*}M_{i}^{j}$, where * represents the set of target countries. Figure 8-B illustrates the top 10 destination countries that saw the largest increase in migration of researchers from the target countries. The analysis reveals the presence of Saudi Arabia, Qatar, and the UAE, which could be attributed to their robust economies, geographical proximity to the target countries, and shared language and religion  (?). These nations have particularly attracted scientists from Egypt, Tunisia, and Jordan. The connection between Egypt and Saudi Arabia has grown significantly, with an average of 258 scientists affiliated with Egyptian institutes relocating to Saudi Arabia each year since the Arab Spring. This trend reflects Saudi Arabia’s influential role in fostering research ties through increased funding, scholarly attention, and attracting scientists. Malaysia, as a majoritarily Islamic country, has emerged as a growing destination for researchers and a source of scholarly attention and funding, most probably because of the “Muslim Brotherhood”, a transnational Islamic movement known for promoting political and social reforms based on Islamic principles founded by Egypt  (?).
Researchers from the target countries migrated to Islamic countries at significantly higher rates compared to countries with other religions, both before and after the Arab Spring. This trend intensified following the Arab Spring, with the share increasing notably. While 5.1% of all immigrant scholars migrated to Islamic countries before and 8.5% after the Arab Spring, the corresponding values for migrant scholars from the target countries were substantially higher, at 33.5% before and 46.2% after (see Fig S8 in Supplementary Materials).
[…]

### v3

@@ Results
The contributions of different countries to the rise in foreign scholarly attention show significant geographical disparities (Fig. 5a). Western nations, known for their high scholarly output, account for the largest share of this growth, whereas contributions from the Global South remain relatively limited. The United States leads, with an average increase of (+172.3 papers/year), followed by the United Kingdom (+ 118.3 papers/year). Malaysia and Saudi Arabia also contribute significantly despite their lower scholarly output.
The increase in scholarly attention of sources countries strongly correlates with their total scholarly output during the period (Pearson correlation: 0.684; Spearman correlation: 0.845, both with $p<0.0001$, see Fig. S4 in the Supplementary Materials) and with their average GDP during this period (Pearson correlation: 0.432; Spearman correlation 0.608, both with $p<0.0001$, see Fig. S5 in the Supplementary Materials). These two factors, GDP and scholarly output, possibly explain these disparities.
@@
To understand the effect of the Arab Spring, we examine the changes in the positions of target countries within the rankings of countries that are sources of attention. We show the change in ranks of target countries for the 10 most important sources of attention in Fig. S6 in the Supplementary Material. A one-way signed rank test indicates a statistically significant increase in the rankings of the government overthrow and civil war groups across all source countries, except for Germany, France, and the UK, at the 0.05 significance level. We find that the ranks of Syria, Libya, Tunisia, and Jordan generally increased in the post-Arab Spring period. Egypt’s rank was already relatively high before the Arab Spring and remained relatively stable. This suggests that the shift in scholarly attention toward these target countries is also significant from the perspective of the source countries. Such a trend may reflect an academic response to regional developments and potential changes in funding priorities or broader shifts in research policy.
@@
Because the distribution of foreign attention’s contribution shares ($A_{i}^{*}$) follows a long-tailed pattern, making upward movement in the rankings is more challenging at lower ranks due to the inertia created by established research infrastructures. For example, a country advancing from rank 2 to rank 1 requires significantly more publications than a country moving from rank 147 to rank 146 among the 147 countries worldwide in our work. To account for these disparities, we measure the normalized rank change of country $i$ as $NRC(A_{i}^{*})=\frac{\Delta r(A_{i}^{*})}{N-r(A_{i}^{*})}$, which takes into account rank inertia  (?). Here, $N$ represents the total number of countries (in this case, $N=147$) and $r(A_{i}^{*})$ denotes the ranking of the country $i$ within the global distribution of scholarly attention directed toward target countries. Figure 5b shows that, in relative terms, countries like Malaysia, Iran, and Qatar also saw a large change in their attention toward the target countries. Notably, Saudi Arabia and the UK appear in the top 5 countries that saw large relative as well as absolute changes.
@@
When considering only Arab Spring-related research (as in Fig. 2), we observe that Western countries have also contributed significantly to scholarly attention (see Fig. S7 in the Supplementary Materials). However, Türkiye, Egypt, and Israel, countries within the region and directly or indirectly affected by the Arab Spring, also showed notable research activity. Meanwhile, Malaysia ranks 24th with 32 publications, while Saudi Arabia ranks 30th with only 23. A key reason for this may be that most Arab Spring-related research is concentrated in the social sciences, which is not the main focus of research originating from these two countries (Fig. S8 in Supplementary Materials).
@@
Figure 6: Repartition of scholarly attention towards target countries by discipline and source country: Differences in yearly scholarly attention by country, listing the top 10 countries with the greatest differences in the field of Social sciences (a), Health Sciences (b), Life Sciences (c), and Physical Sciences (d).
@@
Figures 6a–d show the top ten sources of foreign attention towards the region by computing the differences in yearly attention before and after the Arab Spring in the four main disciplines. Saudi Arabia emerges as a significant source of scholarly attention in the health, physical, and life sciences, with a particular focus on Egypt. Within the same disciplines, Iran’s research focuses on Oman. Malaysia stands out as a key contributor to scholarly attention in the social and health sciences, particularly about Yemen. Spain’s research focuses on Morocco and Tunisia, likely influenced by geographic proximity. In the social sciences, Western countries show a strong attention to Egypt and Syria, whereas their engagement in the life sciences is comparatively muted.
@@
To assess the impact of the Arab Spring on patterns of scholarly attention, we employ a difference-in-differences (DiD) statistical framework  (?). This approach, commonly used in econometrics and quantitative social sciences, involves employing observational data to simulate an experimental research design. It entails studying the differential effect of a treatment on a “treatment group” versus a “control group” in a natural experiment. This approach allows us to estimate the average “treatment ” effect of the Arab Spring by comparing changes over time in attention to our target countries, or in the three subgroups (government overthrow, civil war, and governmental changes) with those observed in either the rest of the world or a subset of MENA countries that remained relatively less affected during the same period (control groups). See the Materials and Methods section for a detailed explanation of the DiD approach and its application to our situation.
@@
Log Domestic Attention log GDP per capita, log Rpop. Rest of world
@@
Table 1: Difference in differences estimates of the effect of Arab Spring on scholarly attention. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of the Arab Spring on log attention (total, domestic, foreign). Models control for log GDP, population, and research population, where applicable. Each colored column represents estimates from separate regressions, comparing the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes regressions where the parallel prior trend test fails to reject the null hypothesis of no difference in pre-Arab Spring trends at a 5% significance level.
Table 1 presents the difference-in-differences (DiD) analysis of scholarly attention. The columns labeled $\beta$ report the estimated effect of the 2011 Arab Spring on scholarly attention to the target countries. The analysis reveals that the Arab Spring had a significant positive effect ($p$ << 0.05) on scholarly attention towards affected countries compared to the rest of the world, and controlling for GDP, population, and researchers’ population. At the subgroup level, countries experiencing government overthrows and civil wars exhibit a positive treatment effect on total attention, with the civil war subgroup experiencing a stronger effect. In contrast, the effect for countries undergoing governmental changes is only statistically significant when compared to the rest of the MENA region, and it is negative. Further disaggregation by the source of attention offers additional insights. The effect on domestic attention is negative in the civil war subgroup compared to the rest of the world and the MENA region. In contrast, the effect is positive in the government overthrow subgroup, possibly reflecting the role of reform strategies. Regarding foreign attention, the effects are positive and statistically significant in the government overthrow and civil war subgroups, with the civil war subgroup exhibiting a notably higher effect. Conversely, the effect on foreign attention is negative in the governmental changes subgroup compared to the rest of the world and the MENA region.
Collectively, these findings provide statistical evidence that the Arab Spring had an impact on scholarly attention toward the target countries and underscore the heterogeneity of this impact across the subgroups. While political instability, via government overthrows and civil wars, has sparked increased academic interest, governmental changes may dampen this engagement.
@@
Figure 7: Shifts in funding patterns before and after the Arab Spring. (a) Changes in funding before and after the Arab Spring per target country. The dashed line, dotted line, and shaded gray area represent the world’s mean, median, and interquartile range. We observe substantial increases for Egypt, Tunisia, Jordan, and Morocco. (b) & (c) Yearly average domestic and foreign funding before and after the Arab Spring per target country. Domestic funding surged for Egypt and the governmental changes (GC) group, particularly Jordan and Kuwait. Foreign funding grows for the government overthrow group (GO), Morocco, and Jordan. (d) Total funding linked with scholarly attention toward the target countries per discipline. Plain and hashed bars indicate values before and after the Arab Spring, respectively. Although the social sciences hold the smallest share of total funding, their allocation shows the largest increase from 12.4% to 17.6%. (e) Change in scholarly attention towards the target countries as a function of changes in funding per source country.
@@
A way for countries to shift scholarly attention is through their allocation of research funding. For each paper in our dataset, we extract the countries providing funding from the article’s acknowledgment section. Similarly to scholarly attention, we model the funding patterns with a directed weighted network. Nodes once more represent countries, and directed edges capture the funding allocated from country $i$ to study country $j$, namely $F^{j}_{i}$ (see Materials and Methods for details). In this way, we measure funding linked to scholarly attention towards the target countries as the number of papers published per year that mention the target countries and have received funding from specific countries.
@@
Figure 7a displays the changes in funding, $\Delta F_{i}^{*}$, received before and after the Arab Spring. The rise in funding for Egypt, Tunisia, Jordan, and Morocco exceeds the median. Domestic funding for studying the government overthrow and governmental changes groups, except Bahrain, has greatly increased (Fig. 7b). The changes for Egypt, Tunisia, and Kuwait are notable, from below the median to above the median. Meanwhile, foreign funding for research on all target countries has grown substantially (Fig. 7c).
Similarly to the patterns observed in scholarly attention, total funding for research directed at these target countries ($F^{*}_{tot}$) increased across all disciplines following the Arab Spring (Fig. 7d). Although the social sciences received the smallest overall share of funding in both periods, their proportion increased from 12.4% to 17.6%. This growth may explain the rise in scholarly attention to the social sciences, which likely aims to assist researchers in interpreting causes, consequences, and long-term societal impacts.
Funding and scholarly attention towards target countries exhibit a strong correlation (Fig. 7e), with a Pearson correlation coefficient of 0.94 and a Spearman rank correlation of 0.90 ($p<0.0001$). The United States leads in funding changes and scholarly attention, followed by Saudi Arabia, France, Spain, and Germany. This ranking suggests that countries with stronger economies tend to have more available funding. However, the Pearson and Spearman correlation coefficients between funding and GDP per capita are only 0.383 and 0.556, respectively ($p<0.001$), indicating only a moderate correlation (see Fig. S9 in Supplementary Materials). Subsequently, while GDP plays a role in funding availability, some countries provide disproportionately higher funding than others. Countries like Iran, Malaysia, and Australia, despite making minor changes in funding, made relatively large changes in their scholarly attention to the target countries.
@@
To assess whether the Arab Spring significantly influenced funding patterns, we applied the same difference-in-differences framework used for scholarly attention (see Materials and Methods), using the log of total, domestic, and foreign funding as outcome variables while controlling for the log of population and GDP per capita. The results, presented in Table S5 of the Supplementary Materials, indicate that, for some regressions, the pre-treatment parallel trends assumption is not satisfied, and the model fit is generally poor. Nonetheless, the remaining regressions that still meet the required assumptions yield results that align with the patterns shown in Figure 7 (see Table S5 in the Supplementary Materials). The Civil War group shows a significant post-Arab Spring positive effect on funding relative to the rest of the world ( $p<0.01$). The effect is also positive on domestic funding for the governmental overthrow group and foreign funding for the civil war group (both significant at $p<0.01$ ). Together, these results underscore a post-uprising shift in both the level and source of funding, varying by political trajectory.
@@
Significant social or political phenomena, such as a revolution or civil war, typically affect a country’s mobility of scholars. These events can create a brain drain, where the number of researchers leaving the country exceeds those entering, or a brain gain, where researchers who had previously left due to political oppression and dissatisfaction with the former government may return as conditions improve. In the former, the net migration rate of researchers increases, and in the latter, it decreases.
@@
In the context of increased scholarly attention, two scenarios can be considered. First, scientists who leave their country due to political upheavals may remain interested in their homeland and continue conducting research related to it. Second, scientists who return to their country after positive political and economic changes may be motivated to contribute to their country’s advancement. This leads to increased domestic knowledge through their intensified research efforts. With this background, we examine whether rising scholarly attention from a given country to the target countries is accompanied by an increase in scholarly migration to that country.
@@
Figure 8: Shifts in migration patterns before and after the Arab Spring (a), Net migration of researchers for target countries before and after the Arab Spring. In all panels, the gray area represents the interquartile range for global trends, with the dashed line indicating the median and the dotted line showing the mean. The inset in panel a (GO) shows the relevant values for Egypt and Tunisia. (b), Changes in the number of migrations from target countries to destination countries before and after the Arab Spring. We show the top 10 destinations. Islamic and Western countries are among the top 10 migration destinations for researchers from target countries, with special bonds between Saudi Arabia and Egypt. (c) Change in scholarly attention towards the target countries as a function of changes in migrations from the target countries.
@@
We track scholarly migration from and to the target countries using the scholarly migration database created using changes in affiliations recorded in Scopus-indexed articles  (?, ?). Figure 8 shows the shifts in migration patterns before and after the Arab Spring. Before the Arab Spring, Egypt consistently attracted researchers, evidenced by positive net migration rates, except in 2008, when the rate nearly reached zero, which could be attributed to unfavorable economic conditions (Fig. 8a). With the onset of the Arab Spring, an increase in the net migration rate of researchers is observed. This may be due to two factors: researchers who had previously left the country due to economic inefficiencies and political turmoil might have returned because of improved conditions. Additionally, Egypt, being a relatively stable country in the region with reputable universities, has attracted researchers from other target countries. However, after 2015, the net migration rate tends to decrease, potentially due to the effects of the Arab Winter, the resurgence of authoritarianism and Islamic extremism in some target countries. Tunisia exhibits a downward trend, transitioning from a country that attracted talent to one experiencing a brain drain. This process began in 2011, and the only subsequent year when the net migration rate was not negative was 2016, a significant year in Tunisia’s political landscape. In 2016, Tunisia experienced widespread protests triggered by high unemployment, particularly among young graduates, and Tunisia’s prominent political party underwent a significant shift. Syria began experiencing a brain drain immediately following the Arab Spring. Meanwhile, patterns for Yemen and Libya are less informative due to their relatively small number of researchers compared to other countries. No clear pattern is observed within the group of countries undergoing governmental changes.
@@
We now model scholarly migration as a directed network, $M$, where nodes represent countries and directed edges, $M_{i}^{j}$, indicate the number of scholars migrating from country $i$ to country $j$. Figure 8b shows the ten destination countries that saw the largest increase in migration of researchers from the target countries. The analysis reveals that Saudi Arabia, Qatar, and the UAE are among the top ten destinations, which could be attributed to their robust economies, geographical proximity to the target countries, and shared language and religion  (?). Post Arab Spring, 162 more scientists affiliated with Egyptian institutes have relocated to Saudi Arabia each year. Malaysia, a predominantly Islamic country, emerges as a growing destination for researchers. Researchers from the target countries migrated to Islamic countries at significantly higher rates compared to countries with other religions, both before and after the Arab Spring. While 5.1% of worldwide immigrant scholars migrated to Islamic countries before and 8.5% after the Arab Spring, the corresponding values for migrant scholars from the target countries were substantially higher, at 33.5% before and 46.2% after (see Fig. S10 in Supplementary Materials).
Western countries, such as the UK, USA, France, and Germany, are also in the top 10 destinations, likely due to their strong economies and leadership in scientific research. The 10 destination countries are classified as HIC (high-income countries) and Upper-Middle-Income Countries (UMIC) in the Worldbank classification. There exists a weak, significant correlation between the change of migration to a country and its average GDP in the period (Spearman Correlation Coefficient: 0.33, $p<0.0001$, see Fig S11 in Supplementary Materials).
A positive correlation exists between changes in scholarly migration from target countries, $\Delta M^{i}_{*}$, and change in scholarly attention towards them, $\Delta A_{i}^{*}$, (Fig. 8c), demonstrated by a Pearson correlation coefficient of 0.505 and a Spearman rank correlation of 0.725, both of which are highly significant ($p<0.0001$). This alignment suggests that migration patterns may partly explain shifts in scholarly attention.